OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. -3 + yi = x + 6i

OpenStudy (anonymous):

i means (1,0) ???

OpenStudy (anonymous):

move everything over to left side... -3 - x + yi - 6i = 0 then -3 - x = 0 and (y - 6)i = 0 solve

OpenStudy (anonymous):

-3 + yi = x + 6i comparing real and imaginary numbers from both sides we find x=-3 and yi=6i i.e. y=6i/i=6 thus x= -3 and y=6 make the equation true